OpenStudy (anonymous):

The volume of a gas varies directly as its temperature measure in degrees Kelvin (K) and inversely as its pressure. When the temperature of a certain gas is 250 K and its pressure is 76.2cm of mercury, its volume is 750 cm^3. To the nearest cubic centimeter, find the volume of this gas when its temperature is 300 K and its pressure is 73.66 cm of mercury.

OpenStudy (anonymous):

V - volume T - temperatyre P - Pressure according to the quesiton: V~T/P where ~ means proportional. so we can write : V = C * (T/P) where C is some constant of the gas. now we can use the data to find C and then using this C we can find the Volume. so first we have: V1 = 750 , T1 = 250 , P = 76.2 plug it into V = C * (T/P) and get : C = 228.6 now in order to find the requested volume we will plug the data : V = 228.6 * 300/73.66 = 931cm3